1) A loan of RM 10,000 at a flat rate of 10% per annum was repaid by making 24 monthly instalment.Find

A) the total interest charged,

B)the monthly payment,and

C)the outstanding loan just after the tenth payment using the rule of 78.
Answer #1

Total interest charged = Principal * Rate * Time = 10,000 × 10% × 2 = RM 2,000

Monthly payment = ( loan + interest) / Total Monthly payments = 12,000/24 = RM 500

Using the Rule of 78, outstanding balance = RN – I(1 + 2 + 3+ ... + N)/( 1 + 2 + 3 + ... + n)

where R = monthly payment

N = number of payments yet to be settled

I = total interest charged

n = total number of payments

= (500 × 14) – 2,000 [(1 + 2 + 3 + ... + 14)/( 1 + 2 + 3 + ... + 24)

= RM 6,300